OpenStudy (chris215):

Write and then solve for the differential equation for the statement: "The rate of change of y with respect to x is inversely proportional to y^2."

OpenStudy (chris215):

dy/dx=k/y^2, y^2*dy=k*dx, y^3/3=kx+C, where 'C' is any constant. The final expression for 'y' as a function of 'x' is y=(3*(kx+C))^(1/3), or y = (3kx+A)^(1/3), where 'A' is any real

OpenStudy (welshfella):

looks good

OpenStudy (chris215):

Thank you!